How many milliliters of water have to be added to  134.1 mL of  0.27 M HCl to reduce the concentration to 0.10 M HCl?


1
Expert's answer
2021-10-25T03:46:59-0400

M1V1= M2V2

134.1×0.27= 0.1×X

36.18= 0.1×

x= 361.8

= 361.8-134.1= 227.7ml